Find the mean variance and standard deviation of the following probability



Z. 2. 6. 8. 10. 12


P(z). 0.35. 0.24. 0.20. 0.15. 0.06

1
Expert's answer
2022-03-08T12:52:11-0500

"\mu=2*0.35+6*0.24+8*0.20+10*0.15+12*0.06=5.96."

"\sigma^2=2^2*0.35+6^2*0.24+8^2*0.20+10^2*0.15+12^2*0.06-5.96^2=10.9584."

"\sigma=\sqrt{10.9584}=3.3103."